OpenStudy (anonymous):

Add or subtract 3/x-3 - 5/x-2

OpenStudy (anonymous):

add or subtract 7/x-3 + 3/x-5

OpenStudy (compassionate):

\[\frac{ 3 }{x-3 } - \frac{ 5 }{ x - 2 }\]

OpenStudy (compassionate):

^ Is that correct?

OpenStudy (anonymous):

Yes.

OpenStudy (compassionate):

\[\frac{ 3(5) }{ (x - 3)5 } - \frac{ 5(3)}{(x - 2)5)}\]

OpenStudy (compassionate):

^ Simplify it.

OpenStudy (anonymous):

i don't know how to though

OpenStudy (compassionate):

Just multiply everything by 5.

OpenStudy (anonymous):

15/x-15 -15/x-10?

OpenStudy (compassionate):

Okay, good. Now you can subtract the numerators and add the denominators. Because the 15 - 15 = 0 You're left with the two denominators. (x - 15)(x - 10) Now solve that.

OpenStudy (anonymous):

am i adding those? x-25?

OpenStudy (compassionate):

x * x x * -10 -15 * x -15 * -10 You have a binomial. You're foiling.

OpenStudy (anonymous):

i don't understand.

OpenStudy (compassionate):

Okay - it's easy. If I was given the terms: 3/x - 5/x I want to find the least common multiple (LCM) The least common multiple of 3 and 5 is 15. So, what do I need to multiply 3 and 5 by to get 15? I need to multiply 3 by 5 and 5 by 3. But what I do to the numerator I must also do to the denominator to remain equality. 15/3x - 15/5x Now the 15 - 15 = 0 So I'm left with: 3x - 5x = -2x Simple, right?

OpenStudy (anonymous):

Oh ok i get it now.

OpenStudy (compassionate):

So, in this case, in the denominator we're left with x - 3 and x - 5 We need to use the FOIL method on these (X- 3)(X - 5) Make sense?

OpenStudy (anonymous):

Yes.

OpenStudy (compassionate):

Okay, so multiple that out. I'll brb. When I get back I'll check your answer.

OpenStudy (anonymous):

Neevermind i just saw that this question was entered in erroor due to the answer choices.

OpenStudy (compassionate):

Okay. :P As always, Best Responses are appreciated.